### v2

# MFAS: Emotion Recognition through Multiple Perspectives Fusion Architecture Search Emulating Human Cognition

Abstract
Speech Emotion Recognition (SER) is an important research topic in human-computer interaction. Many recent works directly extract emotional cues using pre-trained knowledge, often disregarding the appropriateness of it. Taking inspiration from human cognition, we partition speech information into Textual-related Emotional Content (TEC), approximating text, and Speech-related Emotional Content (SEC), encapsulating acoustic information. Given that humans perceive TEC and SEC as discrete and dimensional respectively, we capture them using quantization-based and continuous-based knowledge respectively. Moreover, considering the diverse contextual demands for adjusting TEC using different aspects of SEC as humans do, we propose a new search space for SER and employ neural architecture search for automated design of optimal adjustment strategy. Building on these results, we propose a novel SER framework named Multiple perspective Fusion Architecture Search (MFAS). Experimental results show that MFAS outperforms existing methods.
*Equal contribution
Index Terms—  speech emotion recognition, multiple perspectives, neural architecture search, different modeling methods.
1 Introduction
Due to its significant contribution to human-computer interaction, Speech Emotion Recognition (SER) has received increasing attention. Researchers have been working towards endowing models with the ability to perceive and recognize emotions akin to humans. While some studies have incorporated various prior knowledge to guide the modeling process, the emergence of highly performant pre-trained models has led many recent works to directly leverage pre-trained knowledge for favorable outcomes. Nevertheless, the absence of prior knowledge in modeling could potentially restrict the model’s ability to extract emotional cues from data. In this paper, we employ pre-trained knowledge and simulate human cognitive processes to model emotional cues in speech, and we validate the significance of such biased induction in SER.
Fig. 1: The overall framework of MFAS.
While both text and speech are vital for emotion recognition [1, 2, 3], obtaining accurate transcribed text in real-world scenarios can be challenging. Thus, we broadly partition emotional cues in speech into two types: Textual-related Emotional Content (TEC), which can be extracted using Automated Speech Recognition (ASR) model to approximate text, and Speech-related Emotional Content (SEC), which can be extracted using pre-trained models to emphasize acoustic information. Many existing pre-trained ASR models, grounded in inductive bias towards pronunciation units, introduce quantization operations to enhance the model’s perception of text [4, 5, 6]. After undergoing two stages of pre-training and fine-tuning, they can efficiently construct a constrained set of discrete units. This discrete knowledge aligns well with the cognitive process of human understanding of TEC in speech.
Unlike TEC, during SEC extraction, existing methods lack explicit biased induction guidance and often rely on leveraging knowledge from pre-trained models [7, 8]. However, emotional information perceived by humans is dimensional, continuous [9, 10]. We follow this cognitive principle and compare the performance of quantitative modeling pre-trained knowledge [5] with continuous modeling pre-trained knowledge [11] in the SER fairly. Experimental results show that pre-trained knowledge based on continuous modeling is better suited for capturing SEC.
The proper utilization of multi-perspective information, such as TEC and SEC, is also a matter of concern. People typically perceive emotions based on explicit emotional words in text. However, in certain contexts, this information can be misleading, such as someone might say ”I’m really happy right now” with emphasis when they are angry. In such cases, individuals adjust their perception of emotional words based on acoustic information like intonation or tone, or both. We follow this principle and utilize SEC to adjust the model’s perception of TEC. However, due to the varying significance of multi-level speech information [12], selecting the optimal level of information is imperative. While exhaustive selection can yield answers, it’s a time-consuming and resource-intensive process. Therefore, we propose a novel search space for SER and utilize Neural Architecture Search (NAS) [13] to automatically design adjustment strategy. Experimental results analyze the necessity of this process.
Building on the above findings, we propose a novel SER framework, called Multiple perspective Fusion Architecture Search (MFAS). As shown in Figure 1, MFAS emulates human cognition by employing diverse knowledge perspectives to extract TEC and SEC from speech, and also offers the capability to automatically utilize SEC to adjust biases in TEC. Finally, both features are employed for emotion analysis. The contributions of this paper can be summarized as follows:
  • We validate that knowledge learned by the pre-trained method based on continuous modeling is better suited for capturing SEC than quantization modeling.
  • We propose a novel search space for SER, and utilize NAS to design adjustment strategy for TEC using SEC.
  • We propose a novel SER framework, called MFAS, to simulate the human cognition. Experimental results show that MFAS outperforms existing methods.
Fig. 2: Unified pre-training framework for different modeling methods. Simultaneously, testing the ability of pre-trained knowledge to capture emotional cues.
2 Datasets
Interactive Emotional Dyadic Motion Capture (IEMOCAP) [14] dataset is a widely adopted benchmark for evaluating emotion recognition models. Following prior research, we merge the ”happy” and ”excited” categories into a single ”happy” and focus on four emotions: angry, sad, happy, and neutral. To ensure a robust evaluation of our method, we adopted the 5-fold leave-one-session-out and 10-fold leave-one-speaker-out validation strategies to generate results.
Surrey Audio-Visual Expressed Emotion (SAVEE) [15] dataset consists of recordings from 4 male actors in 7 different emotions. Following prior research, we adopted the 10-fold cross-validation strategies to generate results.
3 Experimental Setup
In this paper, we adopt Co-attention [7] as our Emotion Analysis module. Thus, in data preprocessing, in alignment with it, we segmented each speech into several 3-second-long segments and extracted the corresponding spectrogram and MFCC features for these segments. The final prediction for each speech was determined by aggregating the predictions from all its segmented parts.
Additionally, we employed an SGD optimizer with a learning rate of 5 to optimize the operation weight $\alpha$, as described in Section 5.1. When dealing with MFAS dual-stream inputs, we double the dimension of the encoding hidden features for both spectrograms and MFCCs. These features are evenly distributed to each stream for co-attention calculations.
Table 1: Speech encoder architecture.
CNN Encoder strides 5, 2, 2, 2, 2, 2, 2
kernel width 10, 3, 3, 3, 3, 2, 2
channel 512
Transformer layer 4
embedding dim. 512
FFN dim. 2048
attention heads 8
4 Pre-trained knowledge for SEC extraction
Table 2: The results of emotion analysis using knowledge learned from different modeling approaches.
IEMOCAP SAVEE
Leave-one-sessiong Leave-one-speaker
Modeling
Method
Model
Setting
V A D UA(%) WA(%) V A D UA(%) WA(%) UA(%) WA(%)
Quantization $B$=2 $W$=2 $N$=4 1.02 0.54 0.71 60.3 59.6 1.07 0.60 0.71 62.4 61.9 37.1 43.5
$B$=4 $W$=2 $N$=16 1.38 0.55 0.62 59.7 58.9 1.38 0.54 0.59 63.0 61.9 35.6 42.9
$B$=2 $W$=8 $N$=64 1.13 0.49 0.68 60.6 60.3 1.06 0.44 0.64 63.8 63.3 42.9 47.9
$B$=2 $W$=12 $N$=144 1.01 0.47 0.65 60.6 60.2 0.95 0.40 0.57 63.2 62.3 45.8 49.2
$B$=4 $W$=8 $N$=4096 0.72 0.28 0.46 60.3 60.2 0.70 0.28 0.45 62.6 62.6 47.3 51.3
Continuous $L$=1 0.71 0.28 0.45 60.5 60.3 0.68 0.28 0.44 62.5 62.6 59.1 61.3
$L$=2 0.71 0.28 0.45 61.4 61.3 0.67 0.27 0.44 62.0 62.5 53.6 57.7
$L$=3 0.71 0.28 0.45 60.2 59.9 0.67 0.28 0.45 62.9 62.6 62.8 62.3
$L$=4 0.70 0.28 0.45 60.6 60.6 0.68 0.28 0.44 62.8 62.5 60.6 63.1
4.1 Description of different modeling methods
Currently, speech modeling methods can be broadly categorized into two approaches: quantization-based reconstruction [4, 5, 6] and continuous frame-based reconstruction [16, 11]. Both approaches aim to understand the content of speech by reconstructing the masked portions. The key difference between these approaches lies in their representation space of speech frames. Formally, we represent the speech as $X=[x_{1},...,x_{T}]$, where $T$ is the number of frames. Given a set of indices $Z=[z_{1},...,z_{T^{\prime}}]$ corresponding to the frames that need to be masked, where $0\leq T^{\prime}\leq T$, the frames at the indices in $Z$ are replaced with a trainable code word $c$, resulting in $\tilde{X}=[...,x_{t},...,c,...]$, where $0\leq t<T$. The model then encodes $\tilde{X}$ to obtain its output $\hat{X}=[\hat{x}_{1},...,\hat{x}_{T}]$.
Quantization-based reconstruction methods discretize the representation space of speech frames by clustering them into $N$ centroids, denoted as $Q=[q_{1},...,q_{N}]$. For each frame in $X[Z]$, the corresponding cluster center is used as the label, yielding a label vector $Y=[q_{i},...,q_{j}]$, where $0\leq i,j\leq N$. The objective of the model is to minimize the discrepancy between $\hat{X}[Z]$ and $Y$. On the other hand, continuous frame reconstruction method directly use the original frames at the masked positions as the labels, i.e., $Y=X[Z]$.
4.2 Hyperparameter Configuration
As shown in Figure 2, we use the same model architecture for SER under different modeling constraints to fairly compare them. The specific parameters of the model structure are shown in Table 1. Following previous works [5, 11], we first utilize a 1D-CNN network to encode the speech and obtain $X$. In the quantization modeling approach, we imitate Wav2vec2 [5] and assign $B$ codebooks to all speech frames, each containing $W$ codewords, resulting in a total of $N=W^{B}$ clustering centers. We uniformly set the mask rate to 65%, and the label vector $Y$ is composed of the clustering centers corresponding to $X[Z]$ and the sampled negative examples. The model is trained with BCEloss for reconstruction. On the other hand, in the continuous modeling approach, we imitate Data2vec [11] and use the mean value of the outputs of the last $L$ encoder layers as the $Y$. The model is trained with MSEloss for reconstruction.
4.3 Experimental results
To comprehensively compare the capabilities of various modeling approaches in capturing SEC, we not only test the models’ capability in discrete emotion recognition but also in dimensional emotion analysis. To highlight the differences between regression results, we utilize mean squared error (MSE) as the evaluation metric for dimensional emotion analysis. In Table 2, we report the MSE metrics of valence (V), activation (A), dominance (D), and report unweighted accuracy (UA), weighted accuracy (WA).
As demonstrated by the performance in V, A, and D, in dimensional emotion analysis, knowledge acquired through quantization modeling exhibits notable limitations. The performance significantly improves as the number of quantization units, $N$, increases. However, it is only when the number of $N$ is sufficiently large that this method approaches the performance of the continuous modeling. In the continuous modeling, the variation of the encoding layers $L$ as reconstructed labels does not noticeably affect performance.
For discrete emotion recognition, in the four-class IEMOCAP task, the quantization modeling shows comparable results to the continuous modeling. However, in more complex emotional categorizations, such as the seven-class SAVEE task, quantization modeling leads to a noticeable performance decline. Similarly, inadequate $N$ numbers restrict model performance, while various settings of continuous modeling exhibit relatively consistent performance.
Based on the above experiments, we can conclude that leveraging knowledge acquired through pre-training methods based on continuous modeling enables the capture of more comprehensive emotional information from speech. Whether in discrete emotion recognition or dimensional emotion analysis, the pre-trained knowledge derived from continuous modeling proves to be a superior choice for extracting SEC.
5 Multiple Perspective Fusion Architecture Search
Table 3: Performance comparison with existing works. Evaluation measures are UA / WA for IEMOCAP and UAR / WAR for SAVEE
Model Year
IEMOCAP
Leave-one-session
Model Year
IEMOCAP
Leave-one-speaker
Model Year SAVEE
IS09-classification [17] 2019 63.8 / 68.1 CNN TF Att.pooling [18] 2018 68.1 / 71.8 DCNN [19] 2020    -    / 82.1
CME [20] 2021 73.5 / 72.7 HNSD [21] 2021 72.5 / 70.5 TSP+INCA [22] 2021 83.4 / 84.8
Co-attention [7] 2022 71.1 / 69.8 Co-attention [7] 2022 72.7 / 71.6 CPAC [23] 2022 83.7 / 85.6
Prosody2Vec [24] 2022 73.3 / 72.4 Prosody2Vec [24] 2022 73.9 / 72.7 TIM-Net [25] 2023 77.3 / 79.4
MFAS 2023 74.0 / 71.9 MFAS 2023 74.6 / 73.2 MFAS 2023 86.0 / 86.3
5.1 Description of search space
As illustrated in Figure 1, utilizing an ASR pre-trained model based on quantization modeling allows us to obtain Textual-related Emotional Content (TEC), referred to as $X^{t}$. Similarly, employing a pre-trained model based on continuous modeling with $k$ encoding layers enables us to acquire different levels of Speech-related Emotional Content (SEC), denoted as $[X_{1}^{s},...,X_{m}^{s},...,X_{k}^{s}]$, where $m=\frac{k}{2}$. As the intermediate information generated during speech understanding carries different information [26], we regard $X_{1}^{s}$, $X_{m}^{s}$, and $X_{k}^{s}$ as three levels of information.
  • Due to the constraints of training objective, $X_{k}^{s}$ contains more target-related information.
  • To extract $X_{k}^{s}$ from speech, the intermediate layer need to capture deeper semantic information to fully understand content, denoted as $X_{m}^{s}$.
  • In contrast, the output of the earlier layer, $X_{1}^{s}$, is closer to the raw speech representation.
To prevent interference among different levels of SEC and effectively train the model to adjust TEC using it, during each mini-batch training, we randomly sample one level of SEC, $X^{s}_{i}$, where $i\in[1,m,k]$, for calculation. To ensure that $X^{s}_{1}$ and $X^{s}_{m}$ do not contain target-related information, we detach $X^{s}_{i}$ before using it. We offer a total of eight adjustment operations in the operation pool $\mathbb{O}$, and each operation takes two tensor inputs and yields a tensor output:
$\small\mathrm{Zero}(X^{s}_{i},X^{t})=0$ (1)
$\small\mathrm{Sum}(X^{s}_{i},X^{t})=X^{s}_{i}+X^{t}$ (2)
$\small\mathrm{Attention}(X^{s}_{i},X^{t})=\mathrm{Softmax}(\frac{X^{s}_{i}X^{t}}{\sqrt{C}})X^{t}$ (3)
$\mathrm{ConcatFC}(X^{s}_{i},$ $X^{t})=$ (4)
$\mathrm{ReLU}($ $\mathrm{Linear}(\mathrm{Concat}(X^{s}_{i},X^{t})))$
$\mathrm{ISM}(X^{s}_{i},X^{t})=$ $X^{s}_{i}+\mathrm{Tanh}(\mathrm{Linear}(X^{s}_{i})\ast H)\ast H,$ (5)
$\mathrm{where}\ H=\mathrm{Linear}(X^{t})$
Attention, ConcatFC, and ISM operations have corresponding reverse versions. For example, Attention_r$(X^{s}_{i},X^{t})$ = Attention$(X^{t},X^{s}_{i})$.
Following that, within the Fusion Cell, we employ a differentiable search algorithm [13] to design the adjustment operations. The computational process can be described as:
$X^{t}_{f}=\sum_{o\in O}{\frac{exp(\alpha_{o})}{\sum_{o^{{}^{\prime}}\in O}exp(\alpha_{o^{{}^{\prime}}})}o(X^{s}_{i},X^{t})}$ (6)
where $\alpha\in\mathbb{R}^{|\mathbb{O}|}$ is a trainable weight that measures the importance of different operations, and $|\mathbb{O}|$ denotes the number of operations in $\mathbb{O}$. Both $X^{s}_{k}$ and $X^{t}_{f}$ are subsequently used for emotional analysis as SEC and TEC, respectively.
5.2 Experimental results
5.2.1 Comparison with existing works
Following the conclusions in Section 4, in the experiments of this chapter, we employ the pre-trained knowledge based on continuous modeling from Data2vec to extract SEC from speech. For TEC, we retain the prior knowledge of pronunciation units and capture it using fine-tuned Wav2vec2. Following prior works, we conducted performance tests of MFAS on both the IEMOCAP and SAVEE datasets. As depicted in Table 3, in our experiments on the IEMOCAP dataset, MFAS attains the highest UA score in the Leave-one-session cross-validation strategy, demonstrating competitive WA scores even in the presence of class imbalance. Furthermore, within the Leave-one-speaker cross-validation strategy, MFAS consistently achieves the top UA and WA scores. When applied to the broader range of categories in the SAVEE dataset, MFAS exhibits a significant advantage, surpassing the sota in the 10-fold cross-validation.
Table 4: Performance comparison between different model configurations. Evaluation measures are UA(%) / WA(%).
SEC Encoder TEC Encoder NAS
IEMOCAP
Leave-one-session
Wav2vec2 - 70.6 / 69.6
Data2vec - 72.9 / 71.8
Data2vec Wav2vec2 73.1 / 71.7
Data2vec Wav2vec2 74.0 / 71.9
5.2.2 Comparison with different configurations
We analyze the importance of each module in MFAS by evaluating various model configurations using the Leave-one-session strategy on IEMOCAP. As shown in Table 4, consistent with the findings in Section 4, the pre-trained knowledge based on continuous modeling in Data2vec better captures emotional cues in speech. However, simply incorporating TEC information without adjustment introduces bias, resulting in no significant performance improvement and even a slight decrease in the WA metric. Through the adjustment strategy search by MFAS, the model leverages SEC to mitigate bias in TEC, leading to a further improvement in model performance.
5.2.3 Visualization of Search Results
As shown in Figure 3, in various data partition scenarios, the model automatically selects the optimal level of SEC to adjust TEC. To further demonstrate the performance of MFAS on imbalanced datasets, we present the performance confusion matrix of MFAS compared to the Co-attention method using only Wav2vec2 as the TEC Encoder, as shown in Figure 4.
Fig. 3: The visualization of adjustment strategy search results for the Leave-one-session strategy. Here, $1$, $m$, and $k$ correspond to three levels of information. The red color indicates the selected paths.
(a) Co-attention [7]
(b) MFAS
Fig. 4: Visualization of confusion matrices.
6 Conclusions
In this paper, we validate the significance of knowledge based on continuous modeling in capturing Speech-related Emotional Content (SEC) in speech. Furthermore, we propose a novel search space for adjusting the biases present in Textual-related Emotional Content using SEC. Finally, based on the above results, we introduce a new Speech Emotion Recognition (SER) framework called MFAS. Experimental results show that MFAS outperforms existing methods.